OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (anonymous):

I used the formula cos(theta) = u * v / ||u|| ||v|| to get 38/ sqrt20 * sqrt73 and im not getting any of the listed answers, perhaps im plugging it into the calcu wrong?

OpenStudy (jdoe0001):

hmm what did you get for the dot product?

OpenStudy (anonymous):

38

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

Oh I didnt use cos^-1. Is the answer 6.0deg?

OpenStudy (jdoe0001):

\(\bf \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\to \left(\cfrac{38}{\sqrt{(2)^2+(-4)^2}\cdot \sqrt{(3)^2+(-8)^2}}\right) \\ \quad \\ \to \left(\cfrac{38}{\sqrt{20}\cdot \sqrt{73}}\right)\to \left(\cfrac{38}{\sqrt{1460}}\right)\quad thus\to \theta=cos^{-1}\left(\cfrac{38}{\sqrt{1460}}\right)\)

OpenStudy (jdoe0001):

yeap 6.00900595 degrees, or \(6^o\)

OpenStudy (anonymous):

Thanks, and Im also stuck on this: Evaluate the expression v*w Given the vectors r=<8,1,-6>; v=<6,7,-3); w=<-7,5,2> Does r play any role into this and would -27 be the wrong answer?